Please help me solve. The ages of two sisters are 9 and 12 respectively. In how many years will the product of their ages be 154
Answer by ikleyn(52798)

.

The formal Algebra way is to write this equation


    (9+x)*(12+x) = 154


and find x = "in how many years . . . ", solving this quadratic equation.



As an alternative, you may notice that 154 is a multiple of 11, so 11, probably, will be the age of the younger sister.


If 11, then it will happen in 2 years, when the older sister will be 12+2 = 14.


So, you check 11*14 = 140+14 = 154, and you see that it is correct.


Thus "2 years" is your ANSWER.